OpenStudy (anonymous):

How do I algebraically find the domain of sqrt (1 - x^2)? (i'll re-post as an equation) I know the answer is [-1,1]. But how do I go about solving it algebraically.

OpenStudy (anonymous):

\[\sqrt{1-x^2}\]

Vocaloid (vocaloid):

set the expression underneath the radical greater or equal to 0 \[1-x^{2} \ge 0 \]

OpenStudy (anonymous):

Is what I have below correct? \[1 - x^2 \ge 0 \] \[1 \ge x^2\] \[\sqrt{1}\ge \sqrt{x^2}\] \[\pm1\ge x\]

Vocaloid (vocaloid):

trying to come up with a good way to explain this, sorry \[1 \ge x^{2}\] results in \[-1 \le x \le 1\]

OpenStudy (anonymous):

Thanks

OpenStudy (solomonzelman):

Yes, you are correct. √1=√(x²) 1=|x| x=±1

OpenStudy (mathstudent55):

The radicand must be non-negative, or \(1 - x^2 \ge 0\) To solve this inequality, factor the left side and solve it as an equation: \((1 + x)(1 - x) = 0\) \(1 + x = 0\) or \(1 - x = 0\) \(x = -1\) or \(x = 1\) Now you have 2 points of interest. These two points of interest divide the number line into 3 regions (not counting the points themselves). 

OpenStudy (mathstudent55):

Since the inequality has the symbol \(\ge\) which contains =, the points of interest are part of the solution, so you put closed dots on them.